# Consumption Smoothing in Metropolis: Evidence from the Working-class Households in Prewar Tokyo

### 26 Jan 2024 v3

@@ Abstract
I analyze the risk-coping behaviors among factory worker households in early 20th-century Tokyo. I digitize a unique daily longitudinal survey dataset on household budgets to examine the extent to which consumption is affected by idiosyncratic shocks. I find that while the households were so vulnerable that the shocks impacted their consumption levels, the income elasticity for food consumption is relatively low in the short-run perspective. The result from mechanism analysis suggests that credit purchases played a role in smoothing the short-run food consumption. The event-study analysis using the adverse health shock as the idiosyncratic income shock confirms the robustness of the results. I also find evidence that the misassignment of payday in data aggregation results in a systematic attenuation bias due to measurement error.
Keywords: consumption smoothing; credit purchases; daily budget data; measurement error; risk-coping strategy; risk-sharing
@@ 1 Introduction
The economic theory illustrates the mechanism behind the smoothing of consumption in complete markets (Arrow and Debreu 1954). A series of empirical studies have tested the smoothness of consumption among households in various economies under systematic empirical specification. Evidence shows that markets in developing economies are often imperfect.111 The early representative literature includes, for example, Townsend (1994), Townsend (1995), Ravallion and Chaudhuri (1997), Grimard (1997), and Fafchamps and Lund (2003). Morduch (1990), Townsend (1995), and Dercon (2004) provide a review of the earlier studies. Attanasio and Pistaferri (2016) and Meyer and Sullivan (2017) provide recent discussions. Since risks are not fully insured in the markets, the households are vulnerable to risk, and the idiosyncratic shocks impact the levels of their consumption. Smoothness of consumption is an important factor in accumulating human capital. [^2] For example, using the floods of 1988 in Bangladesh as a shock, Foster (1995) found that children’s growth patterns are influenced by the inefficiency of the credit markets. Rose (1999) investigates the relationship between consumption smoothing and excess losses of girls, revealing that the favorable rainfall shocks in early life increased the survival probabilities of the girls in rural India. Gertler and Gruber (2002) also offer evidence that the health smoothing of consumption against health shocks is fairly imperfect in Indonesia. It follows that consumption smoothing is an essential agenda for mitigating poverty, especially in developing economies.
Historical economies may provide an ideal opportunity to investigate the means and degrees of consumption smoothing under less developed societies. However, the lack of historical data on longitudinal household budgets has hindered research on consumption smoothing. Instead, economic history literature has focused on what type of risk-coping strategies had been used among working-class households during the industrialization.
Kiesling (1996) focuses on the risk-coping strategies among the cotton workers’ households in Lancashire during the recessions due to the Civil War. Time-series data on public relief in six law towns show that the workers had used savings, interfamily transfers, informal charity, and membership in cooperatives and friendly societies as their primary temporary income sources, and used the poor law as the secondary source. [^3] Under a similar empirical setting, Boyer (1997) showed an alternative view that the formal support from the poor law had been used as a main source of temporary income for the unemployed cotton workers. The relative contribution of the sources of temporary incomes from the informal risk-coping strategies was greater among the workers earning higher wages. An important suggestion from both studies for my empirical setting is that the workers could have used mixed sources of risk-coping strategies, particularly during the supply-side macroeconomic shock. Horrell and Oxley (2000) utilized the household-level survey dataset to analyze the risk-coping devices among working-class households in late 19th century Britain. They showed the important roles of the labor supply adjustments among family members in coping with the income losses of the head. A large number of households also took out most insurance, albeit those benefits were insufficient to compensate for the losses. Scott and Walker (2012) provided evidence on what types of temporary income sources were frequently used among interwar British working-class households. Using the household-level budget survey dataset, they found that those households had access to various institutions such as clothing clubs, hire-purchase, contractural savings, and insurance. The households with less earnings were likely to use credit-based institutions like clubs and hire-purchase, whereas those with higher earnings could use the savings institutions. [^4] Another study focuses on the risk-coping institution for the macroeconomic shocks rather than the idiosyncratic shocks. A recent study by Persaud (2019) revealed that the Indian indentureship contracts for South-South migration had worked as a risk-coping device in India. Another strand of literature in financial history provided supply-side (lender-side) evidence on the functions of the micro-financial institutions for the workers. For example, Guinnane (2001) and O’Connell and Reid (2005) investigated the roles of credit cooperatives in 19th-century Germany and check traders in 20th-century England, respectively.
@@
Weights of evidence from these studies have revealed the roles of historical financial institutions and labor supply adjustments in compensating income losses. However, the extent to which those strategies contributed to smooth household consumption has been understudied in the literature. To bridge these gaps in the body of knowledge, this study explores the consumption behaviors in working-class households in the historical metropolis. I digitized the daily longitudinal budget survey on the factory worker households conducted in Tokyo from 1918 to 1920 and employed a systematic econometric method to obtain comparable estimates for income elasticities. The result shows that the full insurance hypothesis is strongly rejected for overall consumption expenditure. The estimated income elasticity is close to that for the rural households in southern India around 1980. This means that the factory worker households suffered certain consumption losses when they faced idiosyncratic adverse income shocks. However, I also found evidence that while expenses on indispensable consumption categories are partially smoothed, expenses on dispensable categories are likely to be cut in response to the idiosyncratic shocks. The results from mechanism analyses show that credit purchases played an important role in mitigating adverse idiosyncratic shocks to smooth workers’ food consumption.
This work has a few potential contributions in a wide range of literature. As described, how working-class households overcame their economic hardships during industrialization has attracted wider attention in economic history. The previous studies have predominantly focused on the temporary income sources used to deal with the risks. Evidence shows that the working-class households had used a variety of risk-coping institutions (Kiesling 1996; Boyer 1997; Horrell and Oxley 2000; Scott and Walker 2012). However, how much these strategies contributed to consumption smoothing among working-class households has yet to be revealed. This paper advances this literature by assessing household behaviors in smoothing consumption.
Relatedly, the methods employed in the previous studies used the cross-sectional variations in income to estimate the elasticities for the temporary income sources. [^5] For example, Horrell and Oxley (2000) “identify households suffering shortfalls in men’s earnings as those where the actual annual earnings fell 20 percent or more below those predicted from an occupational earnings regression” (p. 48). Although Scott and Walker (2012) used the longitudinal budget information (four-week budgets collected at quarterly intervals) to define “household expenditure as “smooth” if it does not deviate by more than one quintile from the mean in any given quarter” (p. 808), their analyses utilized the cross-sectional variations in the pooled cross-sectional sample. Similarly, while James and Suto (2011) predicted the transitory income using the pooled cross-sectional sample, they never utilized the within variations in the panel units. This means that the estimated elasticity indicates the mean tendency that the households with lower (higher) income were more (less) likely to use a specific risk-coping strategy. In other words, the elasticity is derived from the observations across different households with different income levels. However, since risk-coping behavior is the concept of the within (i.e., time-series) variations in a particular household, it is natural to compare the different time cells in the same household when evaluating the roles of the risk-coping strategies. One contribution of this study is to apply the stylized econometric analysis of panel data to systematically investigate the households’ smoothing mechanisms.
@@
A methodological contribution of this study is to reveal a specific generating process of systematic measurement errors in the aggregated income and expenditure series. There were usually two monthly paydays for the factory workers at that time: one was set in the middle and the other at the end of the month. Therefore, the panel cells based on the calendar months, say unadjusted monthly dataset, lead to a substantial mismatch between the timing of earnings and expenses. [^6] If payday is set soon after the initial date in a certain time cell, it does not cause a mismatch in income and expenditure. However, when payday is set just before the initial date in the cell, it leads to measurement errors because the consumption schedule in that cell shall based on the earnings in the previous cell. In other words, this way of matching assumes that the households in a certain cell make their consumption schedule based on “future” earnings. This assumption is unrealistic, particularly in the historical setting. I found that the aggregation ignoring the timing of paydays causes false rejections in most of the consumption subcategories due to the attenuation bias. My finding highlights the importance of reflecting the timing of paydays into the aggregation. This assertion may invoke careful reviews of the existing results based on the aggregate budget dataset, a typical format widely used previously. In the economics literature, several studies highlight the importance of measurement errors in the consumption smoothing analysis (Cochrane 1991; Nelson 1994; Dynarski et al. 1997; Gervais and Klein 2010). This study identified a specific mechanism causing this issue, which may contribute to reducing the attenuation bias in future studies.
@@
Another historical contribution would be that this study adds evidence on the micro-level behaviors of historical consumers during industrialization. Francks (2009) provides a comprehensive view of the long-run consumption history in Japan from the Edo era to the twentieth century. [^7] There are several business history studies analyzing the relationship between consumption and production. A representative work is Gordon (2013), who investigated the expansion of the sewing machine from many aspects, including the production, distribution, and gender social norms in Japan. However, these studies have little interest in the households’ risk-coping behaviors. This current study focuses on short-run consumption behaviors, which is a different analytical view from Francks (2009). [^8] There also be a few studies focusing on the relationships between household chores and consumption behaviors in interwar households (Tanimoto 2011; 2012). However, looking into the extent and how the local society could ensure the households’ consumption provides valuable insights into this literature on consumption history. By using Tokyo City as a case, this study provides the first evidence of the roles of local financial institutions in mitigating the idiosyncratic income shocks to smooth consumption among working-class households after the First World War (WWI).
Relatedly, a case study for a historical metropolis still needs to be explored in the economics literature. Townsend (1994) and Ravallion and Chaudhuri (1997) provide evidence on risk-sharing in rural villages in southern India, and Gertler and Gruber (2000) used a panel household survey on 216 villages in Indonesia in 1991 and 1993. While Townsend (1995) offers the results for Bangkok in Thailand between 1975 and 1990, there is still little evidence for the urban worker households. [^9] Skoufias and Quisumbing (2005) also provide some analyses on the citizens in several countries in the 1990s. A recent study using the experimental strategy by Aguila et al. (2017) covers households in two Mexican towns between 2008 and 2010. Although a recent study by Ogasawara (2024) analyzes the factory workers’ households in Osaka city around 1920, the scale of Tokyo was roughly twice as large as Osaka. [^10] Online Appendix Table A.1 compares the size of the large cities in Japan based on the population census conducted in 1920. In this light, this study is the first to provide the case of a gigantic city, which may offer helpful baseline results for future comparisons.
@@
The rest of this paper is organized as follows. Section 2 overviews the Japanese economy around WWI and reviews Tsukisima, the target area in my analysis. I summarize the financial institutions that were accessible to working-class households. Section 3 introduces the documents used to construct my panel dataset and discusses the sample characteristics. Section 4 summarizes the empirical strategies and preliminary results. Section 5 concludes.
@@ 2.1 Japanese Economy around the First World War
Between 1885 and 1915, real GNP grew annually by 2.6%, while population increased at 1.1%. [^11] This implies that modern economic growth had started as defined by Simon Kuznets (Miyamoto 2008, p. 60). In 1915, the percentage of the employed population in the primary, secondary, and tertiary industries were 63%, 20%, and 18%, respectively. Based on domestic genuine production, however, these industries accounted for just under 30%, over 30%, and approximately 40%, respectively. In this light, Japan had the foundations of an industrial nation by WWI (Miyamoto 2008, p. 64).
Soon after the outbreak of WWI, exports from European countries declined as they were busy producing munitions. Japan increased exports to the Asian market on behalf of those European countries and exported munitions and foodstuffs to the European Allies. This led to increased exports relative to imports and a current account surplus. Furthermore, given the declines in the export capacity of the European countries, import substitution, in which imported products are substituted with domestic products, had progressed, expanding the domestic market (Nakamura and Odaka 1989, p. 10). In particular, there was a rise in the machinery, metal, and chemical sectors. This led to heavy chemical industrialization and increased the working class (Nakamura 1971, p. 128). In addition, a worldwide shortage of ships stimulated the development of the shipbuilding industry, which proliferated as a leading industry. A virtuous cycle was created in which the shipbuilding industry stimulated demand for the machinery and steel industries (Sawai and Tanimoto 2016, p. 250).
@@
Through this process of heavy and chemical industrialization, the total number of enterprises increased. Year-on-year growth in the number of incorporated firms was 17.8% in 1918, 22.3% in 1919, and 13.5% in 1920. These are indeed the top three years of the 34 years from after the Russo-Japanese War to 1940 (Nakamura and Odaka 1989, p. 22). The GNP growth from 1915-19 was 11.5% in nominal terms and 4.4% in real terms, the highest growth throughout the 1910s and 1920s (Figure 1).
This rapid industrialization affected the Japanese economy in many ways. First, due to export-led economic growth, large inflows of gold increased domestic currency balances, causing prices to soar (Nakamura and Odaka 1989, p. 21). As the number of firms increased, especially in the industrial sector, there was a labor shortage in the labor market. Although wages began to rise around 1916 as a response to rising prices, wage increases failed to keep pace with soaring prices. By the end of the war, the number of labor disputes grew sharply, and the labor movement flourished (Nakamura and Odaka 1989, p. 23).
Furthermore, the rise of business accelerated urbanization as the rural population moved to cities in search of urban job opportunities (Nakamura and Odaka 1989, p. 22). The share of large cities (with people of 50,000 or more) as a percentage of the total population was less than 10% in 1889 but increased to nearly 16% in 1920 (Online Appendix Figure A.1a). The growth rate was exceptionally high during the WWI period, indicating that structural changes through the war accelerated the rise of the urban population. Urbanization has increased public investment in large cities, especially in construction investment in social capital such as roads and water systems (Nakamura 1971, p. 142, 147). [^12] The city rents have reportedly increased, making it difficult to find available housing in Tokyo and Osaka (Nakamura and Odaka 1989, p. 23).
Exports in the 1920s were lower than before because of the yen’s appreciation against the real exchange rate due to the accumulation of domestic and foreign species during WWI (Sawai and Tanimoto 2016, p. 253). Since the current account deficit was settled through the disbursement of foreign species owned by the government and the Bank of Japan, there was no reduction in the domestic money supply despite the sizeable current account deficit. This resulted in the domestic prices, especially nontradable goods prices, were kept in high (Sawai and Tanimoto 2016 p. 254). The first half of the 1920s saw a temporary decline in industrial investment as the capital investment boom of the WWI period came to a halt under declining demand. On the other hand, construction investment by the private sector, including electric power and private railways, increased (Nakamura 1971, p.145).
As described, there were short-term business cycles, such as the boom caused by WWI and the subsequent recession. However, taken as a whole, Japan achieved a high economic growth rate relative to other countries in the early 20th century and before WWII (Nakamura 1971, pp. 2-5). Total consumption as a percentage of the GNP remained above 80%. This level is indeed not apparent among other countries (Kuznetz 1962). Importantly, the growth rate of real personal consumption per capita was also high: from 1875 to 1939, the average growth rate of personal consumption was 1.36%, and personal consumption never turned negative except during the Russo-Japanese War period. As shown in Figure 1, the growth rates were particularly high during and immediately after WWI, with average growth rates reaching 112% in 1915-19 and 118% in 1920-24. This was a high level even by international standards (Nakamura 1971, p. 27).
Expanding personal consumption expenditures was essential for developing the domestic market during the economic growth. Behind this, the population living in large cities increased rapidly by nearly 16%. In this light, it is essential to focus on the consumption behavior of urban households when considering the Japanese economy at that time, which achieved a high economic growth rate by international standards. However, due to the low availability of statistical data, progress has yet to be made in analyzing the consumption behavior of urban households (Nakamura 1971, pp. 30-31).

### 25 Feb 2024 v6

@@ Abstract
I analyze the risk-coping behaviors among factory worker households in early 20th-century Tokyo. I digitize a unique daily longitudinal survey dataset on household budgets to examine the extent to which consumption is affected by idiosyncratic shocks. I find that while the households were so vulnerable that the shocks impacted their consumption levels, the income elasticity for food consumption is relatively low in the short-run perspective. The result from mechanism analysis suggests that credit purchases with local retailers played a role in smoothing short-run food consumption. The event-study analysis using the adverse health shock as the idiosyncratic income shock confirms the robustness of the results. I also find evidence that the misassignment of payday in data aggregation results in a systematic attenuation bias due to measurement error in the standard consumption smoothing regressions.
Keywords: consumption smoothing; credit purchases; daily budget data; local retailers; measurement error; risk-coping strategy; risk-sharing
@@ 1 Introduction
Economic theory illustrates the mechanism behind the smoothing of consumption in complete markets (Arrow and Debreu 1954). A series of empirical studies have tested the smoothness of consumption among households in various economies under systematic empirical specification. Evidence shows that markets in contemporary developing countries are often imperfect.111 The early representative literature includes, for example, Townsend (1994), Townsend (1995), Ravallion and Chaudhuri (1997), Grimard (1997), and Fafchamps and Lund (2003). Morduch (1999), Townsend (1995), and Dercon (2004) provide a review of the earlier studies. Attanasio and Pistaferri (2016) and Meyer and Sullivan (2023) provide recent discussions. The households are vulnerable to risk, and the idiosyncratic shocks impact the levels of their consumption because risks are not fully insured in the markets. Therefore, consumption smoothing is an essential agenda for mitigating poverty, especially in developing economies. [^2] Smoothness of consumption is also an important factor in human capital formation. For example, Foster (1995) shows that children’s growth patterns are influenced by the inefficiency of the credit markets using the floods of 1988 in Bangladesh as a shock. Rose (1999) investigates the relationship between consumption smoothing and excess losses of girls, revealing that the favorable rainfall shocks in early life increased the survival probabilities of the girls in rural India. Gertler and Gruber (2002) also offers evidence that the smoothing of consumption against health shocks is fairly imperfect in Indonesia.
@@
Historical economies offer experimental environments to investigate the means and degrees of consumption smoothing under less developed societies. Lessons from historical institutions in developed economies shall provide meaningful insights for today’s developing economies. Despite this advantage, little has been known about historical consumption smoothing behaviors, partially because of the lack of data on longitudinal household budgets. This study is to digitize a unique daily longitudinal budget survey on working-class households conducted in Tokyo around 1919 and employed systematic econometric methods to explore the consumption behaviors among factory worker households in a representative manufacturing area. The result shows that the full insurance hypothesis is strongly rejected for overall consumption expenditure. The estimated income elasticity is close to that for the rural households in southern India around 1980. This means that the factory worker households in historical metropolis suffered certain consumption losses when they faced idiosyncratic adverse income shocks. However, I also found evidence that while expenses on dispensable categories are likely to be cut in response to the idiosyncratic shocks, expenses on indispensable consumption categories are partially smoothed in the short-run. Mechanism analyses provide suggestive evidence that credit purchases in the local economy played an important role in mitigating adverse idiosyncratic shocks.
The first contribution of this study is to offer the first case study of a metropolis in a developing economy. Urbanization is now a general phenomenon in the world, and the shares of urban population in many developing countries, especially in East Asia, are reaching a similar level to the developed countries (Henderson and Turner 2020). This fact raises the need for a deeper understanding of consumption behavior among urban working households in the process of economic development. Despite this, there is still little evidence for the urban worker households in the literature. While Townsend (1995) offers the results for Bangkok in Thailand between 1975 and 1990, a large part of the existing studies focused on the rural economies (Townsend 1994; Ravallion and Chaudhuri 1997; Gertler and Gruber 2002). [^3] Recent field experimental studies also focus on the village economies (Comola and Prina 2023; Macours et al. 2022). A few exceptional studies provide case studies on the middle-sized contemporary economies (Skoufias and Quisumbing 2005; Aguila et al. 2017). In historical context, Ogasawara (2024) analyzes the factory workers’ households in Osaka City around 1920, which was roughly half the size of Tokyo City (Online Appendix Table A.1). This study considers a set of homogeneous households in a representative manufacturing area in Tokyo, which provides a valuable empirical setting for testing the risk-sharing behaviors in the target urban economy. [^4] Focusing on the sample sharing the similar occupational and regional characteristics systematically mitigates bias in estimating income elasticity under the standard risk-sharing regressions due to the heterogeneity in the risk preferences (Schulhofer-Wohl 2011; Mazzocco and Saini 2012). I found suggestive evidence of consumption smoothing using credit transactions with local retailers among working-class households in a historical metropolis.
@@
Macroeconomics literature has investigated the theoretical mechanisms behind the false rejections of the null of the risk-sharing hypothesis that are sometimes observed in developed economies. An essential factor disturbing the valid statistical inference is measurement error in aggregate data (Cochrane 1991; Nelson 1994; Dynarski et al. 1997; Gervais and Klein 2010). This study reveals a specific generating process of systematic measurement errors in the aggregated income and expenditure series in the survey dataset. In my empirical setting, the factory workers had two paydays in a month: one was set in the middle and the other at the end of the month. Hence, the panel dataset based on the calendar months leads to a substantial mismatch between the aggregate income and expenditure series. [^5] If payday is set soon after the initial date in a certain time cell, it does not cause a mismatch in income and expenditure. However, when payday is set just before the initial date in the cell, it leads to measurement errors because the consumption schedule in that cell shall based on the earnings in the previous cell. In other words, this way of matching assumes that the households in a certain cell make their consumption schedule based on future earnings, which is unrealistic. I found that the aggregation ignoring this mechanism causes false rejections in most consumption subcategories due to the attenuation bias. This evidence shows the importance of reflecting the timing of paydays in the aggregation and invokes careful reviews of the existing results from the aggregate budget dataset, a typical format widely used previously.
The second contribution is to add a new analytical perspective to economic history studies. How working-class households overcame their economic hardships during the late 19th to the early 20th century Britain has attracted wider attention. Weights of evidence from these studies have revealed the roles of historical financial institutions and labor supply adjustments in compensating income losses in working-class households. [^6] Horrell and Oxley (2000) found the important roles of the labor supply adjustments among family members in coping with the income losses of the head in working-class households in late 19th-century Britain. Scott and Walker (2012) also found that interwar British working-class households had access to various institutions such as clothing clubs, hire-purchase, contractural savings, and insurance. The households with less earnings were likely to use credit-based institutions like clubs and hire-purchase, whereas those with higher incomes could use the savings institutions. Kiesling (1996) and Boyer (1997) focused on the risk-coping behaviors among the cotton workers households in Lancashire during the recession due to Civil War. Another strand of literature in financial history provided supply-side (lender-side) evidence on the functions of the micro-financial institutions for the workers (O’Connell and Reid 2005). However, there is little evidence of how those risk-coping strategies contributed to smooth household consumption. In addition, the previous studies have utilized the cross-sectional variations in income to estimate the elasticities for the temporary income sources. [^7] For example, Horrell and Oxley (2000) and James and Suto (2011) consider the deviations of actual earnings from the predicted earnings to sort out the shocks. While Scott and Walker (2012) considers the smoothness of expenditure rather than income based on the deviation from the mean expenditure, their analyses employed the cross-sectional variations in the pooled cross-sectional sample. An exceptional study is Saaritsa (2011), which shows the estimated elasticity to temporary income from a fixed effect one-way error component model using a longitudinal household budget survey in 1928 Helsinki. Hence, the estimated elasticity tells the mean tendency that the households with lower (higher) income were more (less) likely to use a specific risk-coping strategy. [^8] In other words, the elasticity is derived from the observations across different households with different income levels. However, since risk-coping is a concept about the dynamic behavior of a specific household, comparing the different time cells is more natural to evaluate the roles of the risk-coping strategies. By using the panel dataset, this study implements the stylized econometric method to systematically investigate the households’ smoothing mechanisms to offer comparable estimates.
@@
Another historical contribution is adding micro-level evidence on consumer behaviors during industrialization. Francks (2009) provides a comprehensive view of the long-run consumption history in Japan from the Edo era to the twentieth century. [^9] There are several business history studies analyzing the relationship between consumption and production. A representative work is Gordon (2013), who investigated the expansion of the sewing machine from many aspects, including the production, distribution, and gender social norms in Japan. However, these studies have little interest in the households’ risk-coping behaviors. A few studies investigated on the relationships between household chores and consumption behaviors in interwar households (Tanimoto 2011, 2012). This study employs a different analytical view by focusing on short-run consumption behaviors.Looking into the extent and how the local society could ensure the households’ consumption would provide valuable insights into the consumption history literature.
@@
The rest of this paper is organized as follows. Section 2 overviews the Japanese economy around the First World War (WWI) and reviews Tsukisima, the target area in my analysis. I also summarize the financial institutions that were accessible to working-class households. Section 3 introduces the documents used to construct my panel dataset and discusses the sample characteristics. Section 4 summarizes the empirical strategies and results. Section 5 overviews the credit purchases in the local economy. Section 6 concludes.
@@ 2.1 Japanese Economy around the First World War
Japan had the foundations of an industrial nation by WWI. Between 1885 and 1915, real GNP grew annually by 2.6%, while population increased at 1.1%. [^10] Given that economic growth is faster than population growth, this implies that modern economic growth had started in this period as defined by Simon Kuznets (Miyamoto 2008, pp. 60; 64). In 1915, the percentage of the employed population in the primary, secondary, and tertiary industries were 63%, 20%, and 18%, respectively. Based on domestic genuine production, these industries accounted for just under 30%, over 30%, and approximately 40%, respectively.
Soon after the outbreak of WWI, exports from European countries declined. Japan increased exports to the Asian market on behalf of those European countries and exported munitions and foodstuffs to the European Allies, leading to a current account surplus. Given the declines in the export capacity of the European countries, import substitution had progressed. This expands the domestic market and leads to a rise in the machinery, metal, and chemical sectors. A worldwide shortage of ships also stimulated the development of shipbuilding, which proliferated as a leading industry. The development of the shipbuilding industry stimulated demand for the machinery and steel industries. Through this process of heavy and chemical industrialization, the total number of enterprises increased. [^11] Descriptions about the process of heavy chemical industrialization are based on Nakamura (1971, p. 128), Nakamura and Odaka (1989, p. 10), and Sawai and Tanimoto (2016, p. 250). Year-on-year growth in the number of incorporated firms was 17.8% in 1918, 22.3% in 1919, and 13.5% in 1920. These are the top three years of the 34 years from after the Russo-Japanese War to 1940 (Nakamura and Odaka 1989, p. 22). Figure 1 shows that the GNP growth from 1915-19 was 11.5% in nominal terms (4.4% in real terms), the highest growth throughout the 1910s and 1920s.
This rapid industrialization impacted the Japanese economy in many ways. Due to export-led economic growth, large inflows of gold increased domestic currency balances, causing prices to soar. Although wages began to rise around 1916 as a response to rising prices and labor demand due to the increased number of firms, wage increases failed to keep pace with soaring prices. By the end of the war, the number of labor disputes grew sharply, and the labor movement flourished. Furthermore, the rise of business accelerated urbanization as the rural population moved to cities in search of job opportunities. The share of large cities (with people of 50,000 or more) as a percentage of the total population was less than 10% in 1889 but increased to nearly 16% in 1920 (Online Appendix Figure A.1a). The growth rate was exceptionally high during the WWI period, indicating that structural changes through the war accelerated the rise of the urban population. Urbanization has increased public investment in large cities, especially in construction investment in social capital such as roads and water supply systems. [^12] Descriptions on the export-led growths, urbanization, and investments are from Nakamura and Odaka (1989, pp. 21–23) and Nakamura (1971, pp. 142; 147).
Exports in the 1920s were lower than before because of the appreciation of the yen against the real exchange rate due to the accumulation of domestic and foreign species during WWI. Since the current account deficit was settled through the disbursement of foreign species owned by the government and the Bank of Japan, there was no reduction in the domestic money supply despite the sizeable current account deficit. This resulted in the domestic prices, especially prices of nontradable goods, being kept high. The first half of the 1920s saw a temporary decline in industrial investment as the capital investment boom of the WWI period came to a halt under declining demand. [^13] On the other hand, construction investment by the private sector, including electric power and private railways, increased (Nakamura 1971, p. 145). Descriptions about the post-war economy in this paragraph are from Sawai and Tanimoto (2016, pp. 253–254).
@@
As described, there were short-term business cycles, such as the boom caused by WWI and the subsequent recession. However, taken as a whole, Japan achieved a high economic growth rate relative to other countries in the early 20th century and before WWII. As shown in Figure 1, the growth rates of personal consumption expenditure were particularly high during and immediately after WWI, with average growth rates reaching 112% in 1915-19 and 118% in 1920-24.141414Total consumption as a percentage of the GNP remained above 80%. Although this level is not apparent among other countries (Kuznets 1962), the growth rate of real personal consumption per capita was relatively high: from 1875 to 1939, the average growth rate of personal consumption was 1.36%, and the rates never turned negative except during the Russo-Japanese War period. This was a high level even by international standards. Given the rapid urbanization, therefore, it is plausible to focus on the consumption behavior of urban households when considering the Japanese economy in this period. However, progress has yet to be made. [^15] Descriptions on the personal consumpsion expenditure are from Nakamura (1971, pp. 2–5; 27; 30–31).